OpenStudy (anonymous):

Solve for x 9^2(x+2) = 27^x-2

OpenStudy (anonymous):

\[\large 9^{2(x+2)}=27^{x-2}\] right?

OpenStudy (anonymous):

since \(9=3^2\) and \(27=3^3\) you can write this as \[\large 3^{6(x+2)}=3^{3(x-2)}\] making \[6(x+2)=3(x-2)\]and that is not too hard to solve

OpenStudy (anonymous):

do we have to make the 6 equal to 3?

OpenStudy (anonymous):

that would be a miracle, wouldn't it?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

oh darn i made a mistake lets try again, this time i go slow and step by step

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

this is where we start \[\large 9^{2(x+2)}=27^{x-2}\] yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then since \(9=3^2\) and \(27=3^3\) we can turn this in to \[\large \left(3^2\right)^{2(x+2)}=\left(3^3\right)^{x-2}\]

OpenStudy (anonymous):

ya true now the bases are the same

OpenStudy (anonymous):

then by the laws of exponents, you multiply the exponents and get \[\large 3^{4(x+2)}=3^{3x-6}\]

OpenStudy (anonymous):

since as you said, the bases are the same, now you know that \[4(x+2)=3x-6\] and that is what you need to solve for \(x\) you good with that?

OpenStudy (anonymous):

just give me a sec to solve it

OpenStudy (anonymous):

k

OpenStudy (anonymous):

4x +2 = 3x-6 1x=-6-2 1x/1 = -8/1 x=-8

OpenStudy (anonymous):

if you start with \[4(x+2)=3x-6\] you have to remove the parentheses on the left using the distributive property first step should be \[4x+8=3x-6\]

OpenStudy (anonymous):

oh ok the 4x-3x = 1x -8-6 = -14

OpenStudy (anonymous):

that is what i get, yes so assuming the first equation was right, that answer should be right \(x=-14\)

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

yw